Astronomy Report – February 27th

Good evening Peter,

Thank you for the precisions concerning the calibration. I will try to use the filters in Photoshop.

Concerning the exoplanet work, on the night from the 26 to 27 of February was an observable transit of TOI 959.01 (from the Tess catalogue) from 2:20 to 6:40 (UTC). The transit depth is supposed to be about 30 ppt. I retraced the light curve of the star during the transit using multi-aperture photometry in AstroImageJ. I was not able to see a transit curve, as you can observe in the attached graph. The curve of TOI has the same trend as the comparison stars (to obtain this graph I had to process the photometry with the 8 first images separately from the following 12 images because the star were displaced in the second part of the image set, the gap you can notice comes from this processing issue).
I don’t know if the predicted transit was wrong, or if the images are too noisy to measure the transit properly (although 30 ppt is quite a big transit depth).

The next transit to image for TOI 959.01 will occur in the night of 3rd March. I hope this night we will have a clear sky. Do you think I need a longer exposure time, or shall I try with the same (80 sec)?

For the Musk Observatory, tomorrow the crew engineer and I will check the screw and see what we can do, I will keep you updated!

I saw in Skynet that the weather at Montana Learning Center will be cloudy tonight, therefore I will not submit any observations today.

Wishing you a nice Sunday evening,


Crew Photos – February 27th

Please find attached the daily photos. The photo of day is the one that portrays the Crew cleaning the Hab.

We apologize for sending 5 pictures instead of 6. Given that it was Sunday, our activities were not as diverse as they usually are;

Thank you very much for your understanding,


Journalist Report – February 27th

Sol 6 – Rest Day, simple day

Today was our first rest day. In fact, almost. We did two short EVAs, one for changing batteries for our outdoor devices, and another one for Mathéo to do his first session with the HoloLens Augmented Reality Experiment.

Otherwise, we spent the day together in the Hab, cleaning our place, cooking, speaking about our lives, and watching movies. Mars is starting to feel like home. We loved this time, and I think these kinds of days will be crucial for future long-term space missions and also our analog one which is filled by activities and experiments: breaks are essential. This first week has been really intense, and everyone here needs this time of rest, with simple things. Have the time to take a step back, to appreciate together, to think… And have a clear mind to begin the upcoming week!

And more, even if I have not spoken much of it the last few days, each time we spend “everyday life moments” together, each Crew member has the same unity feeling. Particularly when we sit all around a table to have lunch, to have a coffee, when we discuss on the sofa… I think our crew is deeply united, benevolent to the others, and this mission might create an indestructible link between us, because of our common work and because here we represent the daily human contact to each other, almost like a family. These moments have thus entirely their part of importance in a crewed mission.

For now, I will end this report here, because my crew mates are waiting for me to begin the movie Interstellar, and I do not want to miss living this moment! See you tomorrow!

EVA Report – February 27th

Crew 263 EVA Report 27-02-2022

EVA # 8

Author: Elena Lopez Contreras Gonzalez

Purpose of EVA: Augmented Reality application testing with a HoloLens device. This AR tool has been designed for astronauts and future geological fieldwork on the Moon and Mars.

Start time: 15:00

End time: 15:15

Narrative: Mathéo (the astronaut wearing the HoloLens) and Elena (the support astronaut) entered the airlock at 14:54h. Inside the airlock, Mathéo launched the application successfully. There was no need for troubleshooting or restarts.

At 15:00h they exited the airlock. The application was functioning nominally. The first waypoint was placed in the RAM. Due to the luminosity, the menu was not as visible at it should have been. Therefore, it took a couple of minutes to find a shadowed place for the menu to be visible. The waypoint was successfully placed reported to HabCom. The pictures were correctly taken as well.

Then, waypoints were placed in the Hab, the GreenHab, the Sciende Dome and the Musk’s Solar Observatory. All were successfully placed by Mathéo and there was no need for troubleshooting or application restarts. However, a second waypoint was visible in the Science Dome. The current hypothesis is that it was placed by the previous astronaut to test the application.

Afterwards, a quick tour of the station was done, to check the waypoint placements. All of them had disappeared, with the exception of the ones placed in the Science Dome. The working hypothesis is that it was too bright outside for the waypoints to be seen –as the Science Dome was shadowed-.

Mathéo and Elena entered the airlock at 15:10h after having successfully completed the two tours around the station. The pressurization was completed at 15:15h, time at which they re-entered the Hab. Mathéo had been support astronaut twice before, therefore, he was already familiar with the application and the potential issues that he could encounter.

EVA went nominally and the application was successfully tested.

EVA Report – February 27th

Crew 263 EVA Report 27-02-2022

EVA # 7

Author: Nicolas WATTELLE

Purpose of EVA: Atmospheric Experiment maintenance (batterie change, MegaAres Setup)

The atmospheric experiment includes 5 devices: LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a Weather station. The aim of this experiment is to do testing of Martian-Atmospheric Study Equipment and their maintenance by human operators. Performing their maintenance and set-up while wearing spacesuits and following EVA protocols is the purpose of the experiment. All devices are located at the same place.

Batteries for the LOAC, MegaAres, the Electrical Field Mill and Purple Air needs to be changed daily. Hence, during this maintenance EVA, the 3 batteries (the Field Mill and MegaAres use the same battery) needed to be replaced.

Start time: 08: 58

End time: 09:55

Narrative: EVA Participants entered the airlock at 08:58. Depressurization ended at 09:03. EVA participants checked the rover batteries just in case. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 84.7 Battery Charge – 100%

Curiosity: Usage Duration – 168.7 Battery Charge – 100%

The EVA participants arrived at the Atmospheric Experiment location at 09:10. The 3 batteries were replaced. Marine was in charge of changing LOAC battery, she did it without any problem to declare. Nicolas changed the Purple Air and Electrical Field Mill batteries, and tried to recover data. But there, while taking out the waterproof protection, condensation appeared inside the box. Then it was impossible to turn it on again. Matheo set the MegaAres antenna, the installation went well, the measurement started correctly.

Nicolas’ radio moved out from his ear at 9:19, but stayed inside the helmet. He could still hear the radio, but had to make efforts speaking to be heard.

At 09:46, the participants left the atmospheric experiment location.

Operations Report – February 27th

Crew 263 Operations Report 27-02-2022

SOL: 6

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Opportunity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Curiosity rover used: No

Hours: 168.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 13 gallons

Water (static tank): 417 gallons

Water (loft tank): 42 gallons

Water Meter: 01561467 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: NA

Summary of suits and radios: The microphone of an earpiece is not working anymore, it would be great to have a new one if possible. Thank you!

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: Shannon worked all day to fix the SOC, the generator is working correctly except that the SOC is not calculating correctly.

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: NA

Sol Summary – February 27th

Crew 263 Sol Summary Report 27-02-2022

Sol: 6

Summary Title: Rest day

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: Today was our rest day, except for two very short EVAs. We thoroughly cleaned the Hab. We watched movies, played games. I worked a bit on my water consumption monitoring experiment. Shannon worked all day to fix the SOC issue, we are very grateful for her work.

Anomalies in work: None.

Weather: Mostly sunny, thin clouds, cold.

Crew Physical Status: Very good.

EVA: Two EVAs, one in the morning to change the batteries of the atmospheric experiments, and one in the afternoon for an application testing with a HoloLens. Both went well.

Reports to be filed: Sol Summary, 2 EVA Requests, 2 EVA Reports, Operations Report, Journalist Report, HSO Report.

Support Requested: None except the spare earpiece mentioned in the Operations report.

Copyright © The Mars Society. All rights reserved. | Main Site